# Picosecond Precision Heavy Ion Detector for Hypernuclei Lifetime Studies

### v2

@@ 2 Heavy Ion Detector Concept and Design
The Radio Frequency Heavy Ion Detector (RF HID) presented here builds on this RFT principle for applications in nuclear physics, specifically for direct lifetime measurements of heavy hypernuclei. Figure 1 illustrates its operational setup. Primary electron, photon, or proton bunches, driven by the RF system of the accelerator, hit the target (1) and produce secondary electrons (SE) along with reaction products such as prompt and delayed fission fragments (FF). Upon exiting the target, the FF, in turn, generate tens of SEs on both sides of the target, which are accelerated by a voltage (2.5 kV) applied between a cathode (target) and an accelerating electrode (2). These electrons are deflected by 90∘ using a permanent magnet (3), collimated (4), focused by an electrostatic lens (5) and circularly scanned by the deflector (6) driven by the RF synthesizer (7) synchronized with the accelerator’s RF system (11). Coordinates of the deflected electrons are measured by the PSD (10).
@@
The PSD, consisting of an MCP array (providing an electron multiplication factor of ) and a DLD40 delay-line anode [20], generates position-sensitive signals with nanosecond-scale rise times. The output signals are processed by a digital acquisition system, consisting of Picoscope [21] and a PC. The fragments are captured in two opposing arms of the detector with a time resolution of approximately 10-30 ps for coincident events, depending on the aperture size of the SE collimator. This precise timing of both FF minimizes interference from background signals and dark noise in microchannel plate detectors and improves time resolution by a factor of .
@@
Secondary electrons (SEs) detected in the system originate from two primary sources. The first comprises SEs emitted directly as a result of the primary beam (electrons, photons, or protons) traversing the metallic target. These prompt SEs are generated on both sides of the target surface. The second source involves SEs produced by charged particles resulting from nuclear reactions in the target caused by primary particles. These reaction products, when exiting the target, also emit SEs as they cross its surface. In this case, the times of the produced electrons are determined by the times of nuclear reactions, which are mostly instantaneous.
@@
However, among the nuclear reaction products are also hypernuclei, which are stopped within the target material. These hypernuclei undergo weak decays after characteristic lifetime of approximately 200 ps, emitting charged particles (pions, protons) that subsequently cross the target surface and produce delayed SEs. These delayed SEs are the key signal for hypernuclear lifetime measurements. Accurate reconstruction of the hypernuclei lifetime requires precise measurement of the time distribution of these delayed SEs.
@@
The main challenge lies in distinguishing delayed SEs associated with hypernuclear decay from the overwhelming background of prompt SEs generated directly by the primary beam or by prompt nuclear reactions. In the case of heavy hypernuclei, this separation is significantly simplified: the decay of a hyperon within such a nucleus induces nuclear fission, producing two fission fragments emitted nearly back-to-back (at 180∘). Each of these fragments generates 100 SEs upon crossing the target surface. This SE yield is roughly two orders of magnitude higher than that of a typical primary beam particle. The ratio of delayed and instantaneous secondary electrons associated with prompt fission of nuclei induced by the primary particles will be determined by the probability of the formation of hypernuclei in nuclear reactions. In the case of interactions of 3 GeV pions and protons with an emulsion, this is about 0.09 percent [22] and 0.5 percent in the case of interactions of 1.5 GeV electrons with Bi [23] (see more details in 4).
@@
By selecting an appropriate position-sensitive detector (PSD), the system can be made selectively insensitive to prompt SEs, which arrive at a fixed time with respect to the beam bunch, and hence strike a specific position on the PSD. A 100 ps length shielding, which corresponds to 18-degree sector (figure 2) for 500 MHz scanning case, could be placed in front of the PSD. This will absorb most of the promptly produced SEs and prevent the MCPs from rapid degradation. The lifetime distribution will then be accurately reconstructed from the detected low rate delayed events.
@@ 5 Expected lifetime measurement performance: Monte-Carlo simulations
Figure 9 (left) illustrates an example result of simulated “toy” experiment (a superposition of prompt and delayed events) with and an event ratio of and a detector resolution of =30 ps ( ps). The "true" lifetime was set at 200 ps. The distribution is fitted with an exponential curve in the range ps. Figure 9 (right) shows the distribution of lifetimes reconstructed from 1000 such “toy” experiments, yielding an uncertainty in the extracted lifetime of 10.7 ps. The same value for resulted in the expected uncertainty of 13.5 ps.
@@ 6 Summary
A new time-resolved heavy ion detector, based on circular radio-frequency scanning of secondary electrons, has been designed, built and tested in the laboratory using a synchronized femtosecond laser. The measured time resolution of the detector is approximately 12 ps. This system is particularly suitable for hypernuclear lifetime studies, enabling the separation of high-rate prompt reaction products from delayed events of interest. Laboratory studies of hot carrier lifetime in graphene, which exhibit a time structure similar to that expected for hypernuclear decays, demonstrate the detector’s capability for such measurements. Additionally, the detection of secondary electrons generated by alpha particles has been successfully verified. Detailed Monte Carlo simulations confirm that the expected statistical uncertainty for direct hypernuclear lifetime measurements with this detector is about 10 ps under realistic conditions. These results indicate that the proposed detector concept is feasible and provides a solid basis for future experiments with electron, photon, or proton beams.

### v3

@@ 2 Heavy Ion Detector Concept and Design
The Radio Frequency Heavy Ion Detector (RF HID) presented here builds on this RFT principle for applications in nuclear physics, specifically for direct lifetime measurements of heavy (A200) hypernuclei. Figure 1 illustrates its operational setup. Primary electron, photon, or proton bunches, driven by the RF system of the accelerator, hit the target (1) and produce secondary electrons (SE). Reaction products from beam particles, such as prompt or delayed fission fragments (FF) can also generate SEs, at a level of tens of SEs per fragment. The SEs emitted from both sides of the target are accelerated by an electrical potential (2.5 kV) applied between a cathode (target) and an accelerating electrode (2). They are deflected by 90∘ using a permanent magnet (3), collimated (4), focused by an electrostatic lens (5) and circularly scanned by the deflector (6) driven by the RF synthesizer (7) synchronized with the accelerator’s RF system (11). Coordinates of the deflected electrons are measured by the PSD (10).
@@
The PSD, consisting of an MCP array (providing an electron multiplication factor of ) and a DLD40 delay-line anode [5], generates position-sensitive signals with nanosecond-scale rise times. The output signals are processed by a digital acquisition system, consisting of Picoscope [4] and a PC. The SEs from the fragments are captured in two opposing arms of the detector with a time resolution of approximately 10-30 ps for coincident events, depending on the aperture size of the SE collimator. This precise timing of both FF minimizes interference from background signals and dark noise in the microchannel plate detectors and improves time resolution by a factor of .
@@
Secondary electrons (SEs) detected in the system are originated from two primary sources. The first comprises SEs emitted directly as a result of the primary beam (electrons, photons, or protons) traversing the metallic target. These prompt SEs are generated on both sides of the target surface. The second source involves SEs produced by charged particles resulting from nuclear reactions in the target caused by primary particles. These reaction products, when exiting the target, also emit SEs as they cross its surface. In this case, the times of the produced electrons are determined by the times of nuclear reactions, which are mostly instantaneous.
@@
However, among the nuclear reaction products are also hypernuclei, which are stopped within the target material. These hypernuclei can undergo weak decays after characteristic lifetime of approximately 200 ps, emitting charged particles (pions, protons, fission fragments) that subsequently cross the target surface and produce delayed SEs. These delayed SEs are the key signal for hypernuclear lifetime measurements. Accurate reconstruction of the hypernuclei lifetime requires precise measurement of the time distribution of these delayed SEs.
Another possible outcome to consider is prompt fission of highly excited hypernuclei. In such events, the initial heavy hypernucleus (A200) may undergo prompt fission, producing one fragment and one hyperfragment with a typical mass of A100. Measurements at Jefferson Lab of proton-, pion-, and kaon-associated photofission of bismuth nuclei in the photon energy range GeV showed that the fission probability in coincidence with kaons was approximately 16, about three times larger than proton- and pion-associated fission and roughly twice the inclusive fission probability (see [29] and references therein). This excess was interpreted as a combination of prompt fission () and formation of bound residual states followed by weak nonmesonic decay leading to delayed fission (). Hyperfragments with A100 can also be formed in kaon-associated prompt fission. However, the fissility of nuclei with A100 is approximately three orders of magnitude lower than that of nuclei with A200 [17]. Consequently, the probability that such hyperfragments undergo delayed fission due to weak nonmesonic decay is very small. Therefore, in the proposed experimental configuration, the contribution of delayed fission from A100 hyperfragments is expected to be negligible compared to that of the heavy A200 hypernuclei of interest.
@@
The main challenge lies in distinguishing delayed secondary electrons (SEs) associated with hypernuclear decay from the overwhelming background of prompt SEs generated directly by the primary beam or by prompt nuclear reactions. In the case of heavy hypernuclei, this separation is significantly simplified. The decay of a hyperon in a heavy nucleus induces nuclear fission, producing two fragments emitted nearly back-to-back (180∘). SE yield from each fragment is orders of magnitude larger than the one from a typical primary beam particle. This large SE yield difference and the back-to-back emission geometry allows to use collimation and coincidence requirement between the two detector arms to provide strong suppression of uncorrelated prompt background (see more details in 4).
@@
By selecting an appropriate position-sensitive detector (PSD), the system can be made selectively insensitive to prompt SEs, which arrive at a fixed time with respect to the beam bunch, and hence strike a specific position on the PSD. A 100 ps length shielding, which corresponds to 18-degree sector (figure 2) for 500 MHz scanning case, could be placed in front of the PSD. This shield will absorb the majority of promptly produced SEs and prevent the MCPs from rapid degradation. The lifetime distribution will then be accurately reconstructed from the detected low rate delayed events that fall outside the shielded phase region. The reference zero-time point can be determined under low beam intensity conditions and subsequently shifted, using a phase shifter, to the center of the shielded sector.
@@ 5 Expected lifetime measurement performance: Monte-Carlo simulations
Figure 9 (left) illustrates an example result of simulated “toy” experiment (a superposition of prompt and delayed events) with and an event ratio of and a detector resolution of =30 ps ( ps). The "true" lifetime was set at 200 ps. The distribution is fitted with an exponential curve in the range ps. Figure 9 (right) shows the distribution of lifetimes reconstructed from 1000 such “toy” experiments resulting in an extracted lifetime of ps and an expected uncertainty of ps. Using the same parameters and yields an extracted lifetime of 200.5 ps and an expected uncertainty of 13.5 ps.
@@ 6 Summary
A new time-resolved heavy ion detector, based on circular radio-frequency scanning of secondary electrons, has been designed, built and tested in the laboratory using a synchronized femtosecond laser. The measured time resolution of the detector is approximately 12 ps. This system can be used for the heavy ion precise timing and is particularly suitable for heavy hypernuclear lifetime studies, enabling the separation of high-rate prompt reaction products from delayed events of interest. Laboratory studies of hot carrier lifetime in graphene, which exhibit a time structure similar to that expected for hypernuclear decays, demonstrate the detector’s capability for such measurements. Additionally, the detection of secondary electrons generated by alpha particles has been successfully verified. Detailed Monte Carlo simulations confirm that the expected statistical uncertainty for direct hypernuclear lifetime measurements with this detector is about 10 ps under realistic conditions. These results indicate that the proposed detector concept is feasible and provides a solid basis for future experiments with electron, photon, or proton beams.